a piece of metal weighing 10.0g is heated to raise its temperature from 25.0C to 76.5C
it is found that the metal absorbed 47.6 cal of heat in the process.

a)calculate the specific heat of the metal.
b)based on your results identify the metal.
1
Expert's answer
2015-11-02T12:05:21-0500
Solution
a) Since
qspent = qgained, metal
we write
(mass) (Δt) (Cp, metal) = qspent
47.6 cal equalls to 199.2 J of heat
Δt = 76.5-25.0 = 51.5 °C
(Cp, metal) = 199.2/(10.0x51.5) = 0.387 J g¯1 °C¯1
b) Using the reference data, it is found that this specific heat corresponds to a Copper (Cu)
Answer: 0.387 J g¯1 °C¯1; Cu